OpenStudy (anonymous):

8-(x-15)=6x+3(x+1) Solve for X.

OpenStudy (mathteacher1729):

How far did you get in this problem before getting stuck?

OpenStudy (anonymous):

i ave no idea how to do it , its on my Eoc Study guide

OpenStudy (anonymous):

have*

OpenStudy (mathteacher1729):

Do you know how to distribute and collect like terms? examples: Distribute the four to both the x and negative three because they are in parenthesis 4(x - 3) = 4x - 4*3 = 4x - 12 Collect like terms (x is the 'term') 3x + 2x = 5x 7x - 12x = -4x 2x + 7x - 1x = 8x etc

OpenStudy (mathteacher1729):

You'll have to distribute and collect like terms on both sides of the equation before doing anything else.

OpenStudy (anonymous):

there are no like terms though?dstibuted as in Distibutive property like multiplying?

OpenStudy (mathteacher1729):

Let's focus on the left and right side of the equal sign first: 8-(x-15)=6x+3(x+1) left side 8-(x-15) this is actually 8-1(x-15) (so you have to distribute the -1)

OpenStudy (anonymous):

8-1x +15?

OpenStudy (mathteacher1729):

8-(x-15) is equal to 8-1x +15 Now what about the +8 and the +15 ? You can add them...

OpenStudy (anonymous):

23-1x?

OpenStudy (mathteacher1729):

Yes. So that is the LEFT side of the equation. 23 - x now the RIGHT side of the equal sign... 6x+3(x+1) Do you see where you can distribute here?

OpenStudy (anonymous):

is it 9x+3?

OpenStudy (mathteacher1729):

9x+3 Exactly! :) Now set those two equal to each other... 23 - x = 9x+3

OpenStudy (anonymous):

so then what ?

OpenStudy (anonymous):

combine like terms?

OpenStudy (mathteacher1729):

The trick now is to "get the variable on one side of the equal sign and the constants on the other side". In this case you should add x to both sides.

OpenStudy (anonymous):

23=10x+3, -3 to both sides then divide by 10, the answer is 2? am i correct?

OpenStudy (anonymous):

O.o